# Structure-Preserving Model Order Reduction for Nonlinear DAE Models of Power Networks

### v1

@@ V-A Empirical Controllability Covariance
where $\mathcal{T}^{c}$ represents the set of excitation direction matrices and it contains $q$ orthogonal excitation matrices, each of size ${n_{u}\times n_{u}}$ while $\mathcal{M}^{c}$ denotes the set of excitation magnitudes and it comprises of $s$ positive real numbers, each representing a different magnitude of excitation to apply along the directions specified in $\mathcal{T}^{c}$. Using the above sets, perturbations in the control input for each time step $k$ can be written as $\bm{u}(k)=c_{m}\bm{T}_{l}^{c}\bm{e}_{i}\mathrm{u}(k)+\bm{u}_{0}(0)$ with $c_{m}$ specifying the magnitude, $\bm{T}_{l}^{c}\bm{e}_{i}$ expressing the direction, and $\mathrm{u}(k)$ representing the temporal shape of the perturbation.
@@ V-C Empirical Balanced Model Synthesis
where ${\bm{f}}\left(\cdot\right)={\bm{f}}\left({\bm{x}_{s}},{\bm{u}}_{s},{\bm{w}}_{s}\right)$. From now on, for the sake of simplicity, with a little abuse of notation, we consider $\bm{x}_{s}=\bm{x}$, $\bm{u}_{s}=\bm{u}$, and $\bm{w}_{s}=\bm{w}$. Similarly, let $\bm{A}=\bm{S}^{-1}_{x}{\bm{A}}\bm{S}_{x}$, $\bm{B}_{u}=\bm{S}^{-1}_{x}{\bm{B}_{u}}\bm{S}_{u}$, $\bm{B}_{w}=\bm{S}^{-1}_{x}{\bm{B}_{w}}\bm{S}_{w}$, and $\bm{S}^{-1}_{x}{\bm{f}}\left({\bm{x}_{s}},{\bm{u}}_{s},{\bm{w}}_{s}\right)={\bm{f}}\left({\bm{x}},{\bm{u}},{\bm{w}}\right)$.
The next step is to perform coordinate transformation such as $\tilde{\bm{x}}=\bm{W}{\bm{x}}$ so that the system is balanced. Notice that, for ODE systems using empirical covariance matrices, computing coordinate transformation matrix $\bm{W}$ is straightforward and well-documented, often involving Cholesky factorization or similar techniques–see [21, 22]. However, for NDAE systems, the presence of algebraic equations (which do not exhibit dynamic behavior and represent static constraints) complicates the application of these techniques. The NDAE system evolves in a subspace defined by the differential equations, while algebraic equations restrict this evolution without contributing to the system’s dynamics [20]. Hence, we propose a two-step approach to design a balanced model. In the first step, we perform balancing for the dynamic variables and then in the second step we perform coordinate transformation for algebraic variables so that they can also be truncated, the details are given in the subsequent sections.
@@ V-E Coordinate Transformation of Algebraic Variables
Now as discussed in the previous section, mapping the nonlinearity using $\bm{W}_{R}$ and $\bm{W}_{L}$ as $\bm{W}_{L}{\bm{f}}\left({\bm{W}_{R}\tilde{\bm{x}}},{\bm{w}},{\bm{u}}\right)$ still depends on the dimension of full state vector $\bm{x}$ and thus can be computationally expensive. Then again, one can use the DEIM-based hyper-reduction approach to measure specific points in the state-space and then efficiently interpolate the nonlinearity around the selected points (as done in Eq. 9 of SP-POD-based MOR technique). Having said that, the overall proposed SP-BPOD-based MOR algorithm for the complete NDAE power system model is summarized in Algorithm 2 presented in Appendix B. It is worth mentioning here that in both the proposed SP-POD and SP-BPOD model reduction techniques, the final reduced order model is guaranteed to be an NDAE. In both proposed techniques, the number of dynamic states in the ROM is determined by the dimensions of $\bm{I}_{dr}$; similarly, the number of algebraic variables is controlled by the dimension of user-defined identity matrix $\bm{I}_{ar}$.
@@ VI Case Studies
To implement the proposed SP-POD-based MOR technique, the snapshot data matrices $\bm{X}_{d}$, $\bm{X}_{d}$, and $\bm{X}_{f}$ need to be computed. This is done by carrying out time domain simulations under step disturbance in overall system load demand as follows. Initially, the system operates under steady-state conditions, meaning load demand is exactly equal to generation and thus there are no transients. Then, right at $t>0$ an abrupt disturbance in load demand is applied to the system as: $\bm{P}^{e}_{d}+\bm{Q}^{e}_{d}=(\bm{I}+\Delta_{d})(\bm{P}^{0}_{d}+\bm{Q}^{0}_{d})$, where $\bm{P}^{0}_{d}$ and $\bm{Q}^{0}_{d}$ represent the initial active/reactive load demand and $\bm{P}^{e}_{d}$ and $\bm{Q}^{e}_{d}$ are their respective values after the disturbances. The parameter $\Delta_{d}$ denotes the severity of the disturbances and we selected it as $0.001$. The overall simulation time period is set to be $20$sec and the system transients dynamic and algebraic states data is saved in matrices $\bm{X}_{d}$, $\bm{X}_{d}$, and $\bm{X}_{f}$. These snapshot data is then used in SP-POD Algorithm 1 to design the corresponding ROM.
@@ VI-A Case Study on IEEE 39-bus Systems
Here, we perform MOR on a modified IEEE 39-bus systems. The 39-bus system consists of 9 conventional power plants (steam and hydro-based), one solar farm at Bus 36, and a motor-based load at Bus 14. The one-line diagram of this system is given in Appendix A. The conventional power plants are modeled via detailed $9^{th}$-order dynamics modeling; generator swing equations, turbine and governor models, and excitation system dynamics. The solar power plant is acting in grid-forming mode and is modeled via $12^{th}$-order dynamical model. Further details about the dynamics of the considered power system model are given in Appendix A. The original full 39-bus system consists of $97$ dynamic states and $120$ algebraic variables, i.e $n_{d}=97$, $n_{a}=120$, and $n=217$.
@@
Similar observations have been carried out for the SP-BPOD-based MOR technique and the dimensions of ROM have been determined. After determining suitable dimensions for reduced models, the corresponding ROMs are constructed using the proposed SP-POD and SP-BPOD-based techniques as given in Algorithms 1 and 2. A comparison with BT-based MOR has also been presented as proposed in [10, 27] to showcase the effectiveness of the proposed methodologies. Notice that BT-MOR is applied to LODE systems as it cannot handle NDAE dynamics.
To access the performance of the proposed MOR techniques time-domain simulations are carried out under transient conditions. The dynamic response of the power system is generated by adding step disturbance in load demand right at the start of the simulation. The disturbance in load demand has been added as discussed in the previous section. Both the FOM and ROMs are simulated under these transient conditions and system responses are recorded. To compare the performance between FOM and ROMs the data generated from ROMs are transformed backed and original state vectors are recovered. The results are given in Fig. 2 and 3. For brevity, a couple of dynamic and algebraic variables are shown, we can see the recovered dynamic states (rotor angle and speed of synchronous generator at Bus $35$ and relative angle of solar plant) are close to the original system responses. Similarly, for the algebraic variables shown in Fig. 3 we can see that the original and recovered states are close and accurate to each other. To further evaluate the performance of the proposed MOR techniques, we also compute the root mean square error (RMSE) between the original and the recovered states. Accordingly, the RMSE value for SP-POD is determined to be $0.038$, while for the SP-BOPD-based ROM, it is $0.0019$. We can see that RMSE values are small and thus the ROMs can accurately approximate the full-order system dynamics. These results are also corroborated by Fig. 2 where the error norm is plotted, we can see that the error is close to zero.
@@ VI-B Case Study on 2000-bus Texas System
We want to point out here that while applying SP-BPOD for this case study, the computation of observability covariance $\bm{G}_{0_{11}}$ can become computationally expensive. This is because to compute $\bm{G}_{0_{11}}$ as discussed earlier, we have to perturb each dynamic state independently, and thus, we need to perform $n_{d}$ number of simulation studies. Which can become computationally expensive for the given Texas system with $n_{d}=3102$. However, this can easily be avoided, as discussed in [28]. The idea is only to perturb the dominant POD modes instead of all the dynamic states to approximate $\bm{G}_{0_{11}}$. Hence, $\bm{G}_{0_{11}}$ for this case study has been designed as follows: First, the state vector $\bm{x}$ is coordinately transformed using SP-POD as $\bm{x}(t)\approx\bm{W}_{R}{\bm{z}}(t)$ and then the dynamic states in $\bm{z}$ (which are only $10$) are systematically perturbed to approximate the observability covariance.
@@
We also present the simulation accuracy for the dynamic and algebraic variables separately for this case study in Tab. II. Again, we can verify that the proposed MOR techniques can simultaneously reduce both dynamic and algebraic variables (from $n_{d}=3102$, $n_{a}=564$, and $n=3666$ to $r_{d}=10$, $r_{a}=30$, and $r=40$ in case of SP-POD and to $r_{d}=8$, $r_{a}=25$, and $r=33$ for SP-BPOD) while maintaining very good accuracy. A comparison with state-of-the-art BT-based MOR has also been provided in Tab. II. Notice that to apply BT the system has first been linearized and then converted to an equivalent ODE as in [10]. From Tab. II we can see that the proposed MOR techniques are superior in terms of accuracy and the size of corresponding ROM as compared to BT-LODE. Notice that the reason BT-LODE is providing poor results is because it is designed based on linearized system dynamics, while the proposed techniques consider the complete nonlinear system. Also, BT-LODE cannot reduce the algebraic variables as it is applicable to ODE dynamics only, while the proposed MOR techniques can simultaneously reduce both dynamic and algebraic variables as presented in Tab. II.
@@ VII Concluding Remarks
In this paper, we propose two MOR approaches, namely the SP-POD and SP-BPOD, to simultaneously reduce both dynamic and algebraic variables of NDAE power system models. The SP-POD offers reducing the system order based on the POD modes in the transient simulation data, while in contrast, the SP-BPOD offers designing ROM via balanced-realization. Because of the diagonal structure of the designed coordinate transformation matrix, the corresponding ROMs from both the proposed methods are guaranteed to be NDAE similar to the original power system model. Thus, the proposed techniques preserve the essential differential-algebraic structure of power system models while allowing a smooth transition from reduced order to the full order dynamic and algebraic state variables.
@@ Appendix A Details of the power system dynamics
Borrowing from [2], here we present comprehensive details about the dynamics of power system model used in this study as follows:

### v2

@@ V-A Empirical Controllability Covariance
where $\mathcal{T}^{c}$ represents the set of excitation direction matrices and it contains $q$ orthogonal excitation matrices, each of size ${n_{u}\times n_{u}}$, the set $\mathcal{M}^{c}$ denotes the set of excitation magnitudes and it comprises of $s$ positive real numbers, each representing a different magnitude of excitation to apply along the directions specified in $\mathcal{T}^{c}$, and the set $\mathcal{E}^{c}$ defines the control input to be excited. Using the above sets, perturbations in the control input for each time step $k$ can be written as $\bm{u}(k)=c_{m}\bm{T}_{l}^{c}\bm{e}_{i}\mathrm{u}(k)+\bm{u}_{0}(0)$ with $c_{m}$ specifying the magnitude, $\bm{T}_{l}^{c}\bm{e}_{i}$ expressing the direction, and $\mathrm{u}(k)$ representing the temporal shape of the perturbation.
@@ V-C Empirical Balanced Model Synthesis
where ${\bm{f}}={\bm{f}}\left({\bm{x}_{s}},{\bm{u}}_{s},{\bm{w}}_{s}\right)$. From now on, for the sake of simplicity, with a little abuse of notation, we consider $\bm{x}_{s}=\bm{x}$, $\bm{u}_{s}=\bm{u}$, and $\bm{w}_{s}=\bm{w}$. Similarly, let $\bm{A}=\bm{S}^{-1}_{x}{\bm{A}}\bm{S}_{x}$, $\bm{B}_{u}=\bm{S}^{-1}_{x}{\bm{B}_{u}}\bm{S}_{u}$, $\bm{B}_{w}=\bm{S}^{-1}_{x}{\bm{B}_{w}}\bm{S}_{w}$, and $\bm{S}^{-1}_{x}{\bm{f}}\left({\bm{x}_{s}},{\bm{u}}_{s},{\bm{w}}_{s}\right)={\bm{f}}\left({\bm{x}},{\bm{u}},{\bm{w}}\right)$.
The next step is to perform coordinate transformation such as $\tilde{\bm{x}}=\bm{W}{\bm{x}}$ so that the system is balanced. Notice that, for ODE systems using empirical covariance matrices, computing coordinate transformation matrix $\bm{W}$ is straightforward and well-documented, often involving Cholesky factorization or similar techniques–see [29, 30]. However, for NDAE systems, the presence of algebraic equations (which do not exhibit dynamic behavior and represent static constraints) complicates the application of these techniques. The NDAE system evolves in a subspace defined by the differential equations, while algebraic equations restrict this evolution without contributing to the system’s dynamics [28]. Hence, motivated by [31] we propose a two-step approach to design a balanced model. In the first step, we perform balancing for the dynamic variables and then in the second step we perform coordinate transformation for algebraic variables so that they can also be truncated, the details are given in the subsequent sections.
@@ V-D Balancing Dynamic Variables
3 Compute matrices $\bm{G}_{c_{11}}$ and $\bm{G}_{o_{11}}$ using (15) and (16)
4 Compute $\bm{T}_{1}$ (20), $\bm{T}_{2}$ (23), $\bm{T}_{3}$ (24), and $\bm{T}_{4}$ (25)
5 Construct $\bm{W}_{d}=\bm{T}_{1}\bm{T}_{2}\bm{T}_{3}\bm{T}_{4}$ as in (26)
7 Select $\bm{W}_{a}=\bm{W}_{gc}$
8 Compute $\bm{W}=\mathrm{blkdiag}(\bm{W}_{d},\;\bm{W}_{a})$ as in (6)
10 Design matrix $\bm{T}=\mathrm{blkdiag}(\bm{T}_{d},\;\bm{T}_{a})$ as in (7)
11 Design $\bm{W}_{L}=\bm{W}\bm{T}$ and $\bm{W}_{R}=\bm{W}_{L}^{-1}$
13 Start greedy algorithm to design $\bm{P}_{M}$
14 Construct rank-$p$ approximating basis
$\bm{W}_{fr}=\begin{bmatrix}\bm{w}_{f_{1}}\;,\bm{w}_{f_{2}}\;,\cdots,\bm{w}_{f_{p}}\end{bmatrix}$
15 Choose the first index: $[\rho$, $i_{1}]$ $=\mathrm{max}(\bm{w}_{f_{1}})$
16 Initialize measurement matrix
$\bm{P}_{M_{1}}=\bm{e}_{i_{1}},\,\bm{W}_{fr}=[\bm{w}_{f_{1}}]$
17 for $j=2:p$ do
18       calculate $\bm{c}$ using, $\bm{P}_{M}^{\top}\bm{W}_{fr}\bm{c}=\bm{P}_{M}\bm{w}_{f_{j}}$
19       compute residual, $\bm{d}=\bm{w}_{f_{j}}-\bm{W}_{fp}\bm{c}$
20       update $\bm{P}_{M}$ and $\bm{W}_{fr}$ as follows:
21       $[\rho,i_{j}]=\mathrm{max}(\bm{d})$
22       $\bm{W}_{fr}=[\bm{W}_{fr},\bm{w}_{f_{j}}],\,\bm{P}_{M}=[\bm{P}_{M},\bm{e}_{i_{j}}]$
Algorithm 2 SP-BPOD for power system NDAE models
@@ V-E Coordinate Transformation of Algebraic Variables
Now as discussed in the previous section, mapping the nonlinearity using $\bm{W}_{R}$ and $\bm{W}_{L}$ as $\bm{W}_{L}{\bm{f}}\left({\bm{W}_{R}\tilde{\bm{x}}},{\bm{w}},{\bm{u}}\right)$ still depends on the dimension of full state vector $\bm{x}$ and thus can be computationally expensive. Then again, one can use the DEIM-based hyper-reduction approach to measure specific points in the state-space and then efficiently interpolate the nonlinearity around the selected points (as done in Eq. 9 of SP-POD-based MOR technique). Having said that, the overall proposed SP-BPOD-based MOR algorithm for the complete NDAE power system model is summarized in Algorithm 2. It is worth mentioning here that in both the proposed SP-POD and SP-BPOD model reduction techniques, the final reduced order model is guaranteed to be an NDAE. In both proposed techniques, the number of dynamic states in the ROM is determined by the dimensions of $\bm{I}_{dr}$; similarly, the number of algebraic variables is controlled by the dimension of user-defined identity matrix $\bm{I}_{ar}$.
@@ VI Case Studies
To implement the proposed SP-POD-based MOR technique, the snapshot data matrices $\bm{X}_{d}$, $\bm{X}_{a}$, and $\bm{X}_{f}$ need to be computed. This is done by carrying out time domain simulations under step disturbance in overall system load demand as follows. Initially, the system operates under steady-state conditions, meaning load demand is exactly equal to generation and thus there are no transients. Then, right at $t>0$ an abrupt disturbance in load demand is applied to the system as: $\bm{P}^{e}_{d}+j\bm{Q}^{e}_{d}=(\bm{I}+\Delta_{d})(\bm{P}^{0}_{d}+j\bm{Q}^{0}_{d})$, where $\bm{P}^{0}_{d}$ and $\bm{Q}^{0}_{d}$ represent the initial active/reactive load demand and $\bm{P}^{e}_{d}$ and $\bm{Q}^{e}_{d}$ are their respective values after the disturbances. The parameter $\Delta_{d}$ denotes the severity of the disturbances. The overall simulation time period is set to be $20$sec and the system transients dynamic and algebraic states data is saved in matrices $\bm{X}_{d}$, $\bm{X}_{a}$, and $\bm{X}_{f}$. These snapshot data is then used in SP-POD Algorithm 1 to design the corresponding ROM.
@@ VI-A Case Study on IEEE 39-bus Systems
Here, we perform MOR on a modified IEEE 39-bus systems. The 39-bus system consists of 9 conventional power plants (steam and hydro-based) and one solar farm. The conventional power plants are modeled via detailed $9^{th}$-order dynamics modeling; generator swing equations, turbine and governor models, and excitation system dynamics. The solar power plant is acting in grid-forming mode and is modeled via $12^{th}$-order dynamical model. Further details about the dynamics of the considered power system model are given in Appendix A. The original full 39-bus system consists of $97$ dynamic states and $120$ algebraic variables, i.e $n_{d}=97$, $n_{a}=120$, and $n=217$.
@@
Similar observations have been carried out for the SP-BPOD-based MOR technique and the dimensions of ROM have been determined to be $r_{d}=8$, $r_{a}=3$, and $r=11$. After determining suitable dimensions for reduced models, the corresponding ROMs are constructed using the proposed SP-POD and SP-BPOD-based techniques as given in Algorithms 1 and 2. A comparison with BT-based MOR has also been presented in Tabs. I and II, as proposed in [13, 35] to showcase the effectiveness of the proposed methodologies. Notice that BT-MOR is applied to LODE systems as it cannot handle NDAE dynamics.
To access the performance of the proposed MOR techniques time-domain simulations are carried out under transient conditions. The dynamic response of the power system is generated by adding step disturbance in load demand right at the start of the simulation. The disturbance in load demand has been added by choosing $\Delta_{d}=0.005$ as discussed in the previous section. Both the full-order model (FOM) and ROMs are simulated under these transient conditions and system responses are recorded. To compare the performance between FOM and ROMs the data generated from ROMs are transformed backed and original state vectors are recovered. The results are given in Fig. 2 and 3. For brevity, a couple of dynamic and algebraic variables are shown, we can see the recovered dynamic states (rotor angle and speed of synchronous generator at Bus $35$ and relative angle of solar plant) are close to the original system responses. Similarly, for the algebraic variables shown in Fig. 3 we can see that the original and recovered states are close and accurate to each other. To further evaluate the performance of the proposed MOR techniques, we also compute the root mean square error (RMSE) between the original and the recovered states. Accordingly, the RMSE value for SP-POD is determined to be $0.038$, while for the SP-BOPD-based ROM, it is $0.0019$. We can see that RMSE values are small and thus the ROMs can accurately approximate the full-order system dynamics. These results are also corroborated by Fig. 2 where the error norm is plotted, we can see that the error is close to zero.
@@
Figure 4: Comparison of FOM and ROM for 39-bus system under fault; solar power plant voltage (top-left), relative angle (top-right), reactive power output (bottom left), and rotor angle of Gen. at Bus 30 (bottom right).
Here, we also asses the performance of the proposed technique under fault. To that end, we added a line to ground fault at $t=4$sec on transmission line $4$-$14$ which is then cleared at $50$ msec and $200$ msec from the near and remote end. The simulation results are illustrated in Fig. 4 and we can see that both proposed techniques still yield accurate outcomes.
@@ VI-B Case Study on 2000-bus Texas System
We want to point out here that while applying SP-BPOD for this case study, the computation of observability covariance $\bm{G}_{0_{11}}$ can become computationally expensive. This is because to compute $\bm{G}_{0_{11}}$ as discussed earlier, we have to perturb each dynamic state independently, and thus, we need to perform $n_{d}$ number of simulation studies. Which can become computationally expensive for the given Texas system with $n_{d}=3102$. However, this can be avoided, as discussed in [36]. The idea is only to perturb the dominant POD modes instead of all the dynamic states to approximate $\bm{G}_{0_{11}}$. Additionally, one can consider bypassing the observability covariance computation altogether, relying on controllability alone to derive the ROM as in [32].
@@
We also present the simulation accuracy for the dynamic and algebraic variables separately for this case study in Tab. II. Again, we can verify that the proposed MOR techniques can simultaneously reduce both dynamic and algebraic variables (from $n_{d}=3102$, $n_{a}=564$, and $n=3666$ to $r_{d}=10$, $r_{a}=30$, and $r=40$ in case of SP-POD and to $r_{d}=11$, $r_{a}=31$, and $r=42$ for SP-BPOD) while maintaining very good accuracy. A comparison with state-of-the-art BT-based MOR has also been provided in Tab. II. Notice that to apply BT the system has first been linearized and then converted to an equivalent ODE as in [13]. From Tab. II we can see that the proposed MOR techniques are superior in terms of accuracy and the size of corresponding ROM as compared to BT-LODE. Notice that the reason BT-LODE is providing poor results is because it is designed based on linearized system dynamics, while the proposed techniques consider the complete nonlinear system. Also, BT-LODE cannot reduce the algebraic variables as it is applicable to ODE dynamics only, while the proposed MOR techniques can simultaneously reduce both dynamic and algebraic variables as presented in Tab. II.
Finally, here we would like to compare and summarize the pros and cons of the two proposed MOR techniques. Notice that, both the proposed SP-POD and SP-BPOD are general input-output-based MOR techniques that can be applied to any NDAE power system model represented in state-space format. Both techniques provide accurate ROMs but have their own advantages and limitations.
SP-POD is effective in capturing the dominant behaviors of a system by focusing on the energy content of the modes. However, it does not directly consider the impact of these modes on the system’s controllability or observability. This can be a critical limitation in control applications, as modes with lower energy levels might still significantly influence the system’s response to controls and its observability. On the other hand, SP-BPOD accounts for the controllability and observability of the modes. However, computing the empirical controllability and observability Gramians (as detailed in Eqs. (15) and (16) can be very challenging, requiring systematic perturbation of control inputs (for the controllability Gramian) and states (for the observability Gramian) with fine-tuning, as explained in Sec. VI. That being said, if the overall aim is to perform realtime scalable feedback control and monitoring, SP-BPOD may be the better option. Otherwise, SP-POD could be chosen due to its lesser computational complexity.
@@ VII Concluding Remarks
In this paper, we propose two MOR approaches, namely the SP-POD and SP-BPOD, to simultaneously reduce both dynamic and algebraic variables of NDAE power system models. The SP-POD offers reducing the system order based on the POD modes in the transient simulation data, while in contrast, the SP-BPOD offers designing ROM via balanced-realization. Because of the diagonal structure of the designed coordinate transformation matrix, the corresponding ROMs from both the proposed methods are guaranteed to be NDAE similar to the original power system model. Thus, the proposed techniques preserve the essential differential-algebraic structure of power system models while allowing a smooth transition from reduced order to the full order dynamic and algebraic state variables. Regarding the limitations of the proposed work we want to mention here that both the presented MOR techniques operate in a data-driven manner and require system transient data of all the state variables from time-domain simulations. Thus, to maintain a refined and accurate reduced model representation it needs to be updated with every power flow or optimal power flow cycle. Also, in this study, we only considered the dynamics of PV power plants while the models of wind and other renewable resources are neglected. Furthermore, in the proposed SP-BPOD method, the sets $\mathcal{M}^{c}$ and $\mathcal{M}^{0}$, along with the scalars $\alpha_{u}$ and $\alpha_{x}$, can be considered hyper-parameters that need to be tuned according to the specific test system and operator expertise, as there is no systematic method provided for selecting their values.
@@ Appendix A Details of the power system dynamics
Borrowing from [2], here we present comprehensive details about the dynamics of power system model used in this paper. In particular, we present the dynamics of conventional power plants, dynamics of grid-forming solar power plants, load dynamics, and the algebraic power flow equations.